OpenStudy (anonymous):

how do you write y=4x-5 in standard form

OpenStudy (anonymous):

y - mx+b so this means y = 4x + (-5)

OpenStudy (chaise):

I assume you mean point-slope formula. Point slope formula is: y = mx +b. Your equation is already in this standard form.

OpenStudy (a_clan):

ax+by+c=0 is the standard form

OpenStudy (anonymous):

so would it be 4x-y=5

OpenStudy (a_clan):

all variables and constants on the left hand side

OpenStudy (a_clan):

and right hand side is zero

OpenStudy (a_clan):

4x -y-5=0

OpenStudy (anonymous):

so i add 5 to both sides

OpenStudy (a_clan):

4x-y=5 is also the standard form

OpenStudy (anonymous):

ok..so my slope is 4?

OpenStudy (a_clan):

yes

OpenStudy (anonymous):

ok so i know my slope is 4 my y intercept is 4 and my x intercept is -5 so i know i have 2 points that i can graph my x and y...so how do i find the other points

OpenStudy (a_clan):

no y-intercept is -5 and x-intercept is 1/4

OpenStudy (a_clan):

for y-intercept, you have to convert your equation in the form y=mx+c for x-intercept, you have to convert your equation in the form x=my+c

OpenStudy (a_clan):

and once you have x and y intercept, you don't need two points to plot your graph.

OpenStudy (anonymous):

so if i have x and y those are the only points i need

OpenStudy (a_clan):